OpenStudy (anonymous):

1) Simplify: (3x − 5) + (5x + 1) 2) Simplify the following expression: (5x2 + 3x + 4) − (2x2 + 5x − 1). If the final answer is written in the form Ax2 + Bx + C, what is the value of A? Numerical Answers Expected!

OpenStudy (anonymous):

1) (5x2 + 3x + 4) − (2x2 + 5x − 1) = 5x2 + 3x + 4 − 2x2 - 5x + 1 = 5x2 - 2x2 + 3x - 5x + 4 + 1 = 3x2 - 2x + 5. A = 3, B = -2, C = 5

OpenStudy (anonymous):

hellooo.....?

OpenStudy (anonymous):

do u need anymore help or should i leave now